# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v1

@@ 3. Our Method: TraCS
This section presents our methods, TraCS-D and TraCS-C, along with their theoretical analyses and extensions.
3.1. Location Space
@@
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $[a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{sta}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$.  This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces (Cunningham et al., 2021b; Du et al., 2023; Zhang et al., 2023; Jiang et al., 2022).
@@
The core idea behind having various LDP methods for the location space $\mathcal{S}$ stems from its multiple decompositions. One decomposition means $\mathcal{S}$ can be represented by several subspaces. For example, besides the aforementioned longitude-latitude coordinate representation, each location $\tau_{i}\in\mathcal{S}$ can also be represented by a direction component and a distance component. Based on these decompositions, we propose two LDP methods: TraCS-D, which perturbs the direction and distance subspaces, and TraCS-C, which perturbs subspaces of the Cartesian coordinates.
3.2. TraCS-D
@@
Figure 1. Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$.
[…]
Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$.
Direction space $\mathcal{D}_{\varphi}$. In the space $\mathcal{S}\subset$ $\mathbb{R}^{2}$, the direction is represented by the angle $\varphi$ relative to a reference direction. We define the reference direction as the latitude lines, i.e. the $0^{\circ}$ direction. Thus, the direction space $\mathcal{D}_{\varphi}$ is a circular domain $[0,2\pi)$ for any location. Fixing a location $\tau_{i}$, any other location $\tau_{j}\neq\tau_{i}$ has a unique direction with respect to $\tau_{i}$.
Distance space $\mathcal{D}_{r(\varphi)}$. The other subspace, i.e. the distance space $\mathcal{D}_{r(\varphi)}$, is determined by the direction $\varphi$. Its size is the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ in the direction $\varphi$, which is a function of $\varphi$. For the rectangular $\mathcal{S}$, $r(\varphi)$ has a closed-form expression, which is detailed later. Thus, fixing $\tau_{i}$ and a direction $\varphi$, any other location $\tau_{j}$ has a unique distance $r(\varphi)$ with respect to $\tau_{i}$.
By this decomposition, we can represent any location $\tau_{j}\in S$ as coordinates $(\varphi,r(\varphi))\in(\mathcal{D}_{\varphi},\mathcal{D}_{r(\varphi)})$ relative to $\tau_{i}$. As an example, $\tau_{j}$ in Figure 1(a) is represented by $(\varphi_{1},r(\varphi_{1}))$. This representation is unique and reversible: any pair of coordinates $(\varphi,r(\varphi))$ can be mapped back to a unique location $\tau_{j}\in\mathcal{S}$. Thus, if $\tau_{j}$ is a private location, an LDP mechanism applied to $\tau_{j}=(\varphi,r(\varphi))$ is essentially applied to $\varphi$ and $r(\varphi)$. By Composition Theorem 2.2, we can design perturbation mechanisms for $\varphi$ and $r(\varphi)$, respectively, to achieve LDP for $\tau_{j}$.
@@
Figure 2. Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right.
[…]
Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$.
3.2.1. Direction Perturbation
@@
  • The circular domain $[0,2\pi)$ is bounded but not finite, making the unbounded mechanisms and discrete mechanisms inapplicable.
  • The circular domain has a unique distance metric, e.g. the distance between 00 and $2\pi$ is 00, not $2\pi$, making the Euclidean distance-based mechanisms inapplicable.
Specifically, classical mechanisms like Laplace and Gaussian (Dwork and Roth, 2014) add unbounded noise, resulting in an output domain of (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). This output domain makes them inapplicable to the circular domain $[0,2\pi)$. Additionally, these mechanisms are designed for the distance metric $d(y,x)=|y-x|$, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as the $k$-RR mechanism (Kairouz et al., 2014), require a finite domain size $|k|$, making them unsuitable for direct application to $[0,2\pi)$.
To illustrate the necessity and superiority of TraCS-D’s piecewise-based design, we start with a strawman approach that extends the $k$-RR mechanism to the circular domain $[0,2\pi)$.
Strawman approach: $k$-RR + uniform sampling. ATP (Zhang et al., 2023) divides the direction space $[0,2\pi)$ into $k$ sectors and applies the $k$-RR mechanism. This approach guarantees LDP for the $k$ sectors but not for the entire $[0,2\pi)$. Even so, we can extend this approach to $[0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-RR.
Specifically, the circular domain $[0,2\pi)$ is divided into $k$ sectors $[i-1,i)\cdot 2\pi/k$ for $i=1,2,\ldots,k$. Assume the private direction $\varphi$ falls into the $j$-th sector. Applying the $k$-RR mechanism outputs sector $i=j$ with probability $p=\exp(\varepsilon)/(k-1+\varepsilon)$, or outputs sector $i\neq j$ with probability $p/\exp(\varepsilon)$. Then, we uniformly sample a direction $\varphi^{\prime}$ from the $i$-th sector, i.e. $\varphi^{\prime}\sim U(i-1,i)\cdot 2\pi/k$.
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within each pre-defined sector. If $k$ is small, uniform sampling within each sector leads to significant inner-sector errors, even for large privacy parameter $\varepsilon$. Conversely, too large $k$ values impair the utility of the $k$-RR mechanism (Wang et al., 2017).
@@
We leverage the idea of piecewise-based mechanisms’ design. Specifically, (i) we instantiate a piecewise-based mechanism for the circular domain $[0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, centered around the private direction with high probability and only dependent on the privacy parameter $\varepsilon$.
Definition 3.1 (Direction Perturbation Mechanism).
Given a private direction $\varphi$ and a privacy parameter $\varepsilon$, TraCS-D’s direction perturbation mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ is defined by:
@@
where $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$ is the sampling probability, and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
@@
The above mechanism $\mathcal{M}_{\circ}$ is defined by a piecewise probability distribution with three pieces, where the central piece $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ has a higher probability density $p_{\varepsilon}$. Obviously, this mechanism satisfies LDP for the circular domain $[0,2\pi)$.
A key advantage of TraCS-D’s direction perturbation mechanism is the “dominant” sector. We refer to the central piece $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ as the dominant sector because it centers around the private direction $\varphi$ with high probability. The dominant sector varies with the privacy parameter $\varepsilon$; a larger $\varepsilon$ results in a narrower dominant sector, closer to the private direction $\varphi$, eliminating the inner-sector errors of the strawman approach.
Example 3.2.
Figure 2(b) compares the strawman approach and TraCS-D. Assume the next location $\tau_{i+1}$ is a private location needing direction perturbation, thus the private direction is $\varphi:\tau_{i}^{\prime}\to\tau_{i+1}$. The dominant sector of TraCS-D is much smaller as $\varepsilon$ increases, e.g. $\varepsilon=6$ in Figure 2(b).  Specifically, when $\varepsilon=6$ and $\varphi=\pi/6$, it can be calculated that $p_{\varepsilon}\approx 3.20$ and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})\approx[0.119\pi,0.214\pi)$. Thus, the dominant sector has a size of $0.095\pi\approx 17^{\circ}$ with probability $0.095\pi\times 3.20\approx 0.95$ of being chosen. In comparison, the strawman approach has a dominant sector of $2\pi/6=60^{\circ}$ with probability $0.98$ of being chosen.
Theoretical comparison between the strawman approach and Definition 3.1 is provided in Appendix B.2. This Appendix will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value.
3.2.2. Distance Perturbation
@@
Given location $\tau_{i}$, any other location $\tau_{j}\in\mathcal{S}$ can be represented by a pair of coordinates $(\varphi,r(\varphi))$ relative to $\tau_{i}$. Since we focus on $\mathcal{S}$ as a rectangular area $[a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ has a closed-form expression depending on the direction $\varphi$. Specifically, this distance falls into one of the following four cases:
(1) $|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$
depending on the value of $\varphi$. Appendix B.3 provides the detailed equation form of Equation (1). Figure 3(b) illustrates the first two cases of $|\mathcal{D}_{r(\varphi)}|$. When $\varphi$ falls into $[0,\varphi_{1})$, i.e. Figure 3(a), $|\mathcal{D}_{r(\varphi)}|=(a_{\text{end}}-a_{i})/\cos(\varphi)$. Figure 3(b) shows the second case, i.e. when $\varphi\in[\varphi_{1},\varphi_{2})$, $|\mathcal{D}_{r(\varphi)}|=(b_{\text{end}}-b_{i})/\sin(\varphi)$.
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3. Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
[…]
Two decomposition of location space $S$ at $\tau_{i}$.
With the known distance space $[0,|\mathcal{D}_{r(\varphi)}|]$ and private distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$.111 We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 3.3 and to align with the circular domain. These two domains are equivalent in implementation. Then we design a piecewise-based mechanism for the normalized distance $\overline{r}(\varphi)$.
Definition 3.3 (Distance Perturbation Mechanism).
Given a private distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, TraCS-D’s distance perturbation mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ is defined by:
@@
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property (Dwork and Roth, 2014) ensures it preserves the same privacy level.
Input: Location space $\mathcal{S}$, private trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$

### v2

@@ 3 Our Method: TraCS
This section presents our methods, TraCS-D and TraCS-C, along with their theoretical analyses and extensions. Notations used in this section can be found in Table 3.
3.1 Location Space
@@
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{sta}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$. This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces [9, 10, 37, 16, 29].
@@
The core idea behind having various LDP methods for the location space $\mathcal{S}$ stems from its multiple decompositions. One decomposition means $\mathcal{S}$ can be represented by several subspaces. For example, in addition to the aforementioned longitude-latitude coordinate representation, each location $\tau_{i}\in\mathcal{S}$ can also be represented by a direction component and a distance component. Based on these decompositions, we propose two LDP methods: TraCS-D, which perturbs the direction and distance subspaces, and TraCS-C, which perturbs subspaces of the Cartesian coordinates.
3.2 TraCS-D
@@
Figure 1: Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$. Any other location $\tau_{j}\in\mathcal{S}$ can be represented by direction-distance coordinates $(\varphi,r(\varphi))$ or Cartesian coordinates $(a,b)$.
Direction space $\mathcal{D}_{\varphi}$. In the space $\mathcal{S}\subset$ $\mathbb{R}^{2}$, the direction is represented by the angle $\varphi$ relative to a reference direction. We define the reference direction as the latitude lines, i.e. the $0^{\circ}$ direction. Thus, the direction space $\mathcal{D}_{\varphi}$ is a circular domain $0,2\pi)$ for any location. Fixing a location $\tau_{i}$, any other location $\tau_{j}\neq\tau_{i}$ has a unique direction with respect to $\tau_{i}$.
Distance space $\mathcal{D}_{r(\varphi)}$. The other subspace, i.e. the distance space $\mathcal{D}_{r(\varphi)}$, is determined by the direction $\varphi$. Its size is the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ in the direction $\varphi$, which is a function of $\varphi$. Since $\mathcal{S}$ is rectangular, $r(\varphi)$ has a closed-form expression, which is detailed later. Thus, fixing $\tau_{i}$ and a direction $\varphi$, any other location $\tau_{j}$ has a unique distance $r(\varphi)$ with respect to $\tau_{i}$.
By this decomposition, we can represent any location $\tau_{j}\in S$ as coordinates $(\varphi,r(\varphi))\in(\mathcal{D}_{\varphi},\mathcal{D}_{r(\varphi)})$ relative to $\tau_{i}$. As an example, $\tau_{j}$ in Figure 1(a) is represented by $(\varphi_{1},r(\varphi_{1}))$. This representation is unique and reversible: any pair of coordinates $(\varphi,r(\varphi))$ can be mapped back to a unique location $\tau_{j}\in\mathcal{S}$. Thus, if $\tau_{j}$ is a sensitive location, an LDP mechanism applied to $\tau_{j}=(\varphi,r(\varphi))$ is essentially applied to $\varphi$ and $r(\varphi)$. By Composition Theorem 1, we can design perturbation mechanisms for $\varphi$ and $r(\varphi)$, respectively, to achieve LDP for $\tau_{j}$.
@@
Figure 2: Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right. The dominant sector of TraCS-D narrows as $\varepsilon$ increases, leading to a smaller inner-sector (sampling) error. In contrast, the strawman approach has a fixed dominant sector, which is independent of $\varepsilon$ and leads to a large inner-sector error.
3.2.1 Direction Perturbation
@@
  • The circular domain $0,2\pi)$ is bounded but not finite, making the unbounded mechanisms and discrete mechanisms inapplicable.
  • The circular domain has a unique distance metric, e.g. the distance between 0 and $2\pi$ is 0, not $2\pi$, making the Euclidean distance-based mechanisms inapplicable.
Specifically, classical mechanisms like Laplace and Gaussian [13] add unbounded noise, resulting in an output domain of $(-\infty,+\infty)$. This output domain makes them inapplicable to the circular domain $0,2\pi)$. Additionally, these mechanisms (also existing piecewise-based mechanisms) are designed for the distance metric d(y,x)=|yx|d(y,x)=|y-x|, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as the $k$-$R$ mechanism [18], require a finite domain size $k$, making them unsuitable for direct application to the continuous circular domain $0,2\pi)$.
To highlight the necessity and superiority of TraCS-D’s piecewise-based design, we start by presenting a strawman approach that extends the $k$-$R$ mechanism to the circular domain $0,2\pi)$.
Strawman approach: $k$-$R$ + uniform sampling. A former work ATP [37] divides the direction space $0,2\pi)$ into $k$ sectors and applies the $k$-$R$ mechanism. This approach ensures LDP for the $k$ sectors but not for the entire $0,2\pi)$ domain. Even so, we can extend this approach to $0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-$R$.
Specifically, the circular domain $0,2\pi)$ is divided into $k$ sectors $i-1,i)\cdot 2\pi/k$ for $i=1,2,\ldots,k$. Assume the sensitive direction $\varphi$ falls into the $j$-th sector. Applying the $k$-$R$ mechanism outputs sector $i=j$ with probability $p=\exp(\varepsilon)/(k-1+\exp(\varepsilon))$, or outputs sector $i\neq j$ with probability $p/\exp(\varepsilon)$. Then, we uniformly sample a direction $\varphi^{\prime}$ from the $i$-th sector, i.e. $\varphi^{\prime}\sim U(i-1,i)\cdot 2\pi/k$.
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within a specific sector. That is, even though the perturbed direction $\varphi^{\prime}$ falls into the same sector as the sensitive direction $\varphi$, the distance between $\varphi$ and $\varphi^{\prime}$ may still be large due to the large sector size. This issue is particularly prominent when $k$ is small, as each sector then spans $2\pi/k$ radians. Uniform sampling within such wide sectors introduces substantial inner-sector errors, which persist regardless of how large the privacy parameter $\varepsilon$ is set. On the other hand, increasing $k$ to reduce sector width impairs the utility of the $k$-$R$ mechanism [32].
@@
We leverage the idea of piecewise-based mechanisms’ design. Specifically, (i) we instantiate a piecewise-based mechanism for the circular domain $0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, which is centered around the sensitive direction with high probability and depends solely on the privacy parameter $\varepsilon$, thus eliminating the need for fixed sectors.
Definition 5 (Direction Perturbation Mechanism).
Given a sensitive direction $\varphi$ and a privacy parameter $\varepsilon$, TraCS-D’s direction perturbation mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ is defined by:
@@
where $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$ is the sampling probability, and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
@@
The above mechanism $\mathcal{M}_{\circ}$ is defined by a piecewise probability distribution with three pieces, where the central piece $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ has a higher probability density $p_{\varepsilon}$. As a piecewise-based mechanism, it evidently satisfies LDP for the circular domain $0,2\pi)$; see Appendix B.1 for the proof.
A key advantage of TraCS-D’s direction perturbation mechanism is the “dominant” sector. We refer to the central piece $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ as the dominant sector because it centers around the sensitive direction $\varphi$ with high probability. The dominant sector adapts dynamically to the privacy parameter $\varepsilon$: as $\varepsilon$ increases, the dominant sector becomes narrower and more tightly centered around the sensitive direction $\varphi$. This adaptivity effectively mitigates the inner-sector errors inherent in the strawman approach, where the sector width is fixed and independent of $\varepsilon$.
Example 1.
Figure 2 compares the strawman approach and TraCS-D. Assume the next location $\tau_{i+1}$ is a sensitive location needing direction perturbation, thus the sensitive direction is $\varphi:\tau_{i}^{\prime}\to\tau_{i+1}$. The dominant sector of TraCS-D is much smaller as $\varepsilon$ increases. For instance, when $\varepsilon=6$ and $\varphi=\pi/6$, it can be calculated that $p_{\varepsilon}\approx 3.20$ and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})\approx[0.119\pi,0.214\pi)$. Thus, the dominant sector has a size of $0.095\pi\approx 17^{\circ}$ with probability $0.095\pi\times 3.20\approx 0.95$ of being chosen. In comparison, the strawman approach has a dominant sector of $2\pi/6=60^{\circ}$ with probability $0.98$ of being chosen. With almost the same probability of being chosen, TraCS-D’s dominant sector is significantly narrower than that of the strawman approach, resulting in a much smaller inner-sector (sampling) error.
Detailed comparison between the strawman approach and Definition 5 is provided in Appendix C.2. This Appendix will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value.
3.2.2 Distance Perturbation
@@
Given location $\tau_{i}$, any other location $\tau_{j}\in\mathcal{S}$ can be represented by a pair of coordinates $(\varphi,r(\varphi))$ relative to $\tau_{i}$. Since we focus on $\mathcal{S}$ as a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ has a closed-form expression depending on the direction $\varphi$. Specifically, this distance falls into one of the following four cases:
$|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$ (1)
depending on the value of $\varphi$. Appendix C.3 provides the detailed equation form of Equation (1). Figure 3 illustrates the first two cases of $|\mathcal{D}_{r(\varphi)}|$: when $\varphi$ falls into $0,\varphi_{1})$, i.e. Figure 3(a), $|\mathcal{D}_{r(\varphi)}|=(a_{\text{end}}-a_{i})/\cos(\varphi)$; Figure 3(b) shows the second case, i.e. when $\varphi\in[\varphi_{1},\varphi_{2})$, then $|\mathcal{D}_{r(\varphi)}|=(b_{\text{end}}-b_{i})/\sin(\varphi)$.
With the known distance space $0,|\mathcal{D}_{r(\varphi)}|]$ and sensitive distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$. [^2] We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 6 and to align with the circular domain. These two domains are equivalent in implementation. We then design a piecewise-based mechanism for the normalized distance $\overline{r}(\varphi)$ to guarantee LDP on $[0,1)$.
Definition 6 (Distance Perturbation Mechanism).
Given a sensitive distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, TraCS-D’s distance perturbation mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ is defined by:
@@
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property [13] ensures it preserves the same privacy level.
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3: Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
3.2.3 Workflow of TraCS-D
Combining the direction perturbation and distance perturbation, TraCS-D perturbs each sensitive location in the trajectory.
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-sensitive reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}$ and $\tau_{i+1}$. For each sensitive location $\tau_{i+1}$, it perturbs the direction (red block) and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9). Line 12 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$ in place, which is then used as the reference for the next iteration ($i+1$). Finally, the algorithm outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
Algorithm 1 uses each $\tau_{i}^{\prime}$ as the reference location for the next perturbation. In fact, it can be any non-sensitive location in $\mathcal{S}$. We choose different $\tau_{i}^{\prime}$ for alleviating the impact of a specific reference location.
Input: Rectangular location space $\mathcal{S}$, sensitive trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$